what percentage of 300 is 40?

Answer by srisanj4(2)

Question: What percent of 300 is 40.
Let us say X % of 300 is 40. If I write this in mathematical equation..
X% * 300 = 40
X/100 * 300 = 40
3X = 40
X = 40/3 = 13.33
Hence 13.33 percent of 300 is 40